OpenStudy (anonymous):

PLEASE HELP these are the last to problems on my homework. PLEASE PLEASE HELP!!!!!!!! What is the 8th term of the geometric sequence where a_1 = 1024 and a_3 = 64? What is the sum of a 6–term geometric series if the first term is 8 and the last term is 134,456?

OpenStudy (lgbasallote):

in your first question if a_1 = 1024 and a_3 = 64 then your a_2 = 1024r^2 = 64 square root it... 32r = 8 r = 1/4 so a_8 = a_1r^n-1 a_8 = 1024(1/4)^7 too lazy to do it...but try inputting it in the calculator :DDD

OpenStudy (mertsj):

64=1024r^2 64/1024=r^2 1/4=r \[a _{8}=1024(\frac{1}{4})^7\]

OpenStudy (mertsj):

\[\S _{n}=\frac{a _{1}(1-r^n)}{1-r}=\frac{8(1-7^6)}{1-7}\]

OpenStudy (anonymous):

do any of you know the second question

OpenStudy (mertsj):

Do you even read what people post?

OpenStudy (mertsj):

You ask for help. Someone does the problem and posts it and then you ask if anyone knows how to do the problem.

OpenStudy (lgbasallote):

@Mertsj...why did you do it 8(1-7^6)? your r is 1/4 right?